Circular No. 8936

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

2005 NB_{7}

M. K. Shepard, Bloomsburg University of Pennsylvania; M. C. Nolan, National Astronomy and Ionosphere Center; L. A. M. Benner, J. D. Giorgini, and S. J. Ostro, Jet Propulsion Laboratory, and C. Magri, University of Maine at Farmington, report: “Arecibo delay-Doppler images (2380 MHz, 12.6 cm) obtained on Apr. 11–12 show that the Apollo-type object 2005 NB_{7} (e.g., MPECs 2005-N32, 2005-T82, 2008-F18; MPO 135910) is a binary system. Preliminary estimates of average diameters, based on range estimates at 7.5-m resolution, are 0.5 and 0.2 km (± 0.1) km. The rotation period of the primary has not been determined but, based on the estimated diameter and observed bandwidth, is ≥ 2.5 hr. The semi-major axis of the relative orbit is ≥ 0.6 km; the system orbital period is not known.”

V2468 CYGNI R. J. Rudy, R. W. Russell, and D. K. Lynch, The Aerospace Cor-

poration; and C. E. Woodward, University of Minnesota, report on SpeX observations (wavelength range 0.8–2.5 μm) made with the Infrared Telescope Facility (IRTF) on Mar. 13 and Apr. 12 UT: “V2468 Cyg (cf. IAUCs 8927, 8928) continues to display a very rich emission-line spectrum. Although features of C I, N I, O I, and Fe II still persist, the Ca II infrared triplet, which was very strong in March, is nearly undetectable in the April spectrum. The He I lines are now quite strong, and He II features are beginning to emerge. The O I lines indicate a reddening of E(B–V) = 0.77. There is no indication yet of dust formation.”

V459 VULPECULAE R. W. Russell, D. K. Lynch, R. J. Rudy, The Aerospace Corporation;

and C. E. Woodward, University of Minnesota, report on IRTF SpeX observations, obtained as above on Apr. 12.62 UT: “V459 Vul (cf. IAUC 8907) is one of the small number of classical novae to display C I and coronal lines simultaneously. Of the latter, [S IX] at 1.25 μm is the highest-excitation and [Si VI] at 1.96 μm is the strongest. He II lines and the unidentified novae lines are present as well. The O I lines, which are still quite strong, indicate a reddening of E(B-V) = 1.0, some of which is probably local to the nova since the spectrum exhibits thermal emission from dust beyond 1.5 μm.”

2008 April 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green